Problem 1 is based on the discussion of comparative advantage in Landsburgs textbook. The illustrative test questions contain similar questions (see the questions about food and clothing production in Asia and America). A one line response is sufficient for each of parts A through F. For Part G, provide a one paragraph solution. Only wine and flour are produced in Canaan and Aram: Jacob, living in Canaan, needs 20 hours to produce a barrel of wine and 10 hours to * produce a sack of flour (for bread). Labans sons, living in Aram, need 50 hours to produce a barel of wine and 20 hours to produce a sack of flour Rachel and Leah (Jacobs wives and Labans daughters) are wine and flour traders who travel between Canaan and Aramm A. What is the relative cost of wine (in terms of flour) in Canaan? B. What is the relative cost of wine (in terms of flour) in Aram? C. Where is wine most costly to produce? D. What is the relative cost of flour (in terms of wine) in Canaan? E. What is the relative cost of flour (in terms of wine) in Aram? F. Where is flour most costly to produce? G. Rachel and Leah do one of the following: 1. Carry wine and flour produced in Canaan for sale in Aram 2. Carry wine and flour produced in Aram for sale in Canaan 3. Carry wine produced in Canaan for sale in Aram and flour produced in Aram for sale in Canaan. 4. Carry wine produced in Aram for sale in Canaan and flour produced in Canaan for sale in Aram. Explain which of the above choices Rachel and Leah will do
Answer #1

Solution: Relative cost is basically the opportunity cost of producing another good with the same input.

A. Relative cost of wine in Canaan = 20/10 = 2 sacks of flour

B. Relative cost of wine in Aram = 50/20 = 2.5 sacks of flour

C. Wine is most costly produced in Aram because it takes 50 hours to produce a barrel of wine whereas it takes only 20 hours in Canaan.

D. Relative cost of flour in Canaan = 10/20 = 1/2 barrel of wine

E. Relative cost of flour in Aram = 20/50 = 2/5 barrel of wine

F. Flour is most costly produced in Aram as it takes 20 hours to produce.

G. Since Canaan has absolute advantage in both wine and flour but comparative advantage in wine as compared to Aram, Canaan will produce wine and Aram will produce flour. Thus Rachel and Leah will do option 3.